A transverse wave on a taut string is modeled with the wave function



y(x, t) = Asin(kx-wt)



(0.25m) sin(6.48m-'x- 1.47s1 t)



Find the: (a) amplitude, (b) wavelength, (c) period, (d) direction of



: )



propagation and (e) speed of the wave.

1
Expert's answer
2022-01-12T08:30:09-0500

(a) The amplitude is A=0.25 m.

(b) The wavelength is


"\lambda=\dfrac{2\pi}k=\dfrac{2\pi}{6.48}=0.97\text{ m}."


(c) The period is


"T=\dfrac{2\pi}\omega=4.27\text{ s}."


(d) The direction of propagation is along y-axis.

(e) The speed is


"v=\dfrac \lambda T=0.227\text{ m/s}."